# Towards Automated Identification of Violation Symptoms of Architecture Erosion

### v4

@@ 7.3. Analyzing Software Repositories with Machine Learning and Deep Learning
The application of ML and DL techniques to extract informative knowledge from various artifacts during software development is gaining popularity in the software engineering community. In recent years, a plethora of empirical studies have been conducted to investigate automated techniques for supporting software engineering tasks. Khalajzadeh et al. (Khalajzadeh2022sda, ) conducted an empirical study by extracting and manually analyzing app reviews from 12 OSS projects and developed ML and DL models to automatically detect and classify human-centric issues from app reviews and discussions. Their results show that automated techniques can help developers to recognize and appreciate human-centric issues of end-users more easily. Nasab et al. (Nasab2021ais, ) developed 15 ML and DL models to automatically identify security discussions, and then they collected practitioners’ feedback through a validation survey and confirmed the promising applications of the models in practice.
AlOmar et al. (AlOmar2021SAR, ) proposed an approach to automatically identify and classify self-admitted refactoring in commit messages. They compared their approach with pattern-based and simple random baselines, and the results show that a good performance and a relatively small training dataset is sufficient to classify self-admitted refactoring commits. Hey et al. (Hey2020norbert, ) proposed a novel approach, named NoRBERT, for requirements classification by using the transfer learning capabilities of BERT. They applied their approach to different tasks in the domain of requirements classification, and the results show that their approach improves requirements classification and can be applied to unseen projects with convincing results.
In this work, word embeddings were utilized as word representations and as the input for training ML and DL-based classifiers, as prior studies have demonstrated the effectiveness of word embeddings in capturing rich semantic and syntactic features of words (Ren2019nndSATD, ). We employed ML/DL-related techniques to promote the automated identification of violation symptoms. Although much work has been done on automatic text identification, there is currently no automatic way to identify architecture violations in code review comments. Our approach is not restricted to specific textual information, so future studies can replicate and extend our work with other artifacts.
@@ 8. Conclusions and Future Work
In this work, we attempted to address this issue of automated identification of violation symptoms in code review comments. To this end, we first performed a series of experiments on a dataset of architecture violations from code review comments. Subsequently, to validate the usefulness of our trained ML and DL-based classifiers, we conducted a survey and semi-structured interviews that acquired feedback from the involved developers who discussed architecture violations in code reviews.
@@
Specifically, we developed 15 ML-based and 4 DL-based classifiers to identify violation symptoms from developer discussions in code reviews (i.e., code review comments from four OSS projects in Gerrit). The results show that the SVM classifier based on word2vec pre-trained word embedding performs the best with an F1-score of 0.779. In most cases, classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Furthermore, 200-dimensional pre-trained word embedding models outperform classifiers that use 100 and 300-dimensional models. In addition, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperforms the individual classifiers. Moreover, practitioners’ perception of the usefulness of the classifiers’ results validates the promising applications of automated identification of violation symptoms in practice. Especially, such classifiers can enable and inspire practitioners to find architecturally-related violation issues, prioritize, and eventually handle them.
@@
In the future, we plan to conduct a more in-depth investigation of how practitioners deal with architecture violations in code review, what kind of practices and tools they use in dealing with architecture violations, and how using our approach in their daily job can support the identification of architecture violations. We also plan to further enhance the classifiers by identifying fine-grained types of architecture violations based on a larger dataset of more projects from both open-source and industry projects. In particular, a bigger picture is that we plan to delve into a more in-depth investigation and comparison of the performance of trained classifiers as plug-ins or tools with fine-tuned large language models (such as Llama 2) using real-world scenarios. The goal of such plug-ins or tools for code review platforms is to provide warnings of potential architecture violations to practitioners (e.g., developers and reviewers) for them to re-check the related code changes during code review.
Acknowledgements
We would like to thank all the practitioners who participated in our survey and interviews. This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311 and 62402348.

### v5

@@ 7.3. Analyzing Software Repositories with ML/DL and LLM Techniques
The application of ML/DL and LLM techniques to extract informative knowledge from various artifacts during software development is gaining popularity in the SE community. In recent years, a plethora of empirical studies have been conducted to investigate automated techniques for supporting SE tasks. Khalajzadeh et al. (Khalajzadeh2022sda, ) conducted an empirical study by extracting and manually analyzing app reviews from 12 OSS projects and developed ML/DL models to automatically detect and classify human-centric issues from app reviews and discussions. Their results show that automated techniques can help developers to recognize and appreciate human-centric issues of end-users more easily. Nasab et al. (Nasab2021ais, ) developed 15 ML/DL models to automatically identify security discussions, and then they collected practitioners’ feedback through a validation survey and confirmed the promising applications of the models in practice.
Hou et al. (hou2024large, ) conducted a systematic literature review to explore the applications of LLMs across various phases of software development, from requirements gathering to program repair. Their study provides a comprehensive overview of recent advancements in the applications of LLMs in software engineering tasks, highlighting significant opportunities and critical areas that require further research and development to fully leverage LLMs in software engineering. Ahmed et al. (Ahmed2025can, ) investigated the potential of LLMs to substitute costly human annotation efforts in evaluating SE artifacts. They found that state-of-the-art LLMs can achieve inter-rater agreements comparable to human annotators across various code and code-related tasks. De Martino (Martino2025LLM4MSR, ) et al. presented a mixed-method empirical study combining a rapid literature review and survey to examine the use of LLMs in Mining Software Repositories (MSR). They uncovered 14 approaches and actions to mitigate 9 identified threats in applying LLMs to MSR, and proposed PRIMES 2.0, an enriched six-stage empirical framework with 23 substeps to enhance transparency, reproducibility, and methodological robustness in LLM-based MSR studies.
7.4. Conclusive Summary
Table 11 summarizes the key differences between our work and prior studies. While previous studies addressed relevant topics pertaining to code review, architecture violations, and the use of ML/DL/LLM techniques in mining software repositories, none of them have specifically utilized these automated approaches to classify architecture violations during code review. Building on developers’ perspectives, we further collected feedback from practitioners involved in the code review discussions to evaluate the usefulness of the classifiers. In contrast to previous studies listed in Table 11, our work not only complements prior studies by bridging the gap between identifying violation symptoms and applying automated techniques, but also incorporates practitioners’ feedback to evaluate these automated techniques.
Table 11. Comparison between our work and previous studies
Our Work Previous Studies Comparison
Automatically identifying violation symptoms in code review comments. We explored the application of automated techniques to specifically identify violation-related issues (e.g., violation symptoms) from code review comments. Previous studies (Han2022cmd, ), (ElAsri2019ess, ), (Kashiwa2022SATDmcr, ), (Uchoa2020hdm, ), (Paixao2020bti, ) focused on the code-level issues and typically involved manual review during the code review process. Unlike previous empirical studies that manually analyzed code review comments, we attempted to employ automated techniques to identify architecture-level issues (specifically violation symptoms) and compared the performance differences among different automated techniques.
Investigating architecture violations during code review. Discussions regarding architecture violations can occur in the code review process, and we specifically investigated violation symptoms from textual artifacts like code review comments. Previous studies (Mendoza2021avd, ), (Brunet2012av, ), (Terra2015rsr, ), (Maffort2016mav, ) typically analyzed architecture violations through longitudinal study, tools, and defined rules, and often inspected architecture violations in source code during development or maintenance phases. Compared to previous studies focusing on architecture violations from source code, we proposed an automated approach to identify architecture violations during code review, and our approach can act as a complementary way to modern code review.
Exploring the application of automated techniques to identify violation symptoms. We investigated the feasibility of automated techniques for analyzing violation symptoms by employing various configurations of the trained classifiers. Moreover, we evaluated the usefulness of our approach by collecting practitioners’ feedback. Previous studies (hou2024large, ), (Nasab2021ais, ), (Khalajzadeh2022sda, ), (Ahmed2025can, ), (Martino2025LLM4MSR, ) utilized different automated approaches (such as ML, DL, and LLM techniques) in various SE tasks, including identifying security issues, classifying human-centric app reviews. Differentiated from previous studies utilizing automated techniques to diverse SE tasks, our work specifically explored the feasibility of training traditional ML-based classifiers, DL-based classifiers, as well as state-of-the-art LLM-based classifiers to identify violation symptoms in code review comments. We compared the performance of these models across different configurations to provide more empirical insights. We also solicited practitioners’ viewpoints on the usefulness of our approach.
@@ 8. Conclusions and Future Work
In this work, we attempted to address the issue of automated identification of violation symptoms in code review comments. To this end, we performed a series of experiments on a dataset of architecture violations from code review comments. Subsequently, to validate the usefulness of our trained ML/DL-based classifiers, we conducted a survey and semi-structured interviews that acquired feedback from the involved developers who discussed architecture violations in code reviews. Furthermore, to compare LLM-based classifiers with traditional ML/DL-based classifiers, we performed the same classification task using three state-of-the-art LLMs (i.e., GPT-4o, Qwen-2.5, and DeepSeek-R1).
@@
Specifically, we developed 15 ML-based and 4 DL-based classifiers to identify violation symptoms from developer discussions in code reviews (i.e., code review comments from four OSS projects in Gerrit). The results show that the SVM classifier based on word2vec pre-trained word embedding performs the best with an F1-score of 0.779. In most cases, classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Furthermore, 200-dimensional pre-trained word embedding models outperform classifiers that use 100 and 300-dimensional models. In addition, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperforms the individual classifiers. Moreover, practitioners’ perception of the usefulness of the classifiers’ results validates the promising applications of automated identification of violation symptoms in practice. Especially, such classifiers can enable and inspire practitioners to find architecturally-related violation issues, prioritize, and eventually handle them. Additionally, LLM-based classifiers indeed outperform traditional ML/DL-based classifiers on identifying violation symptoms, and GPT-4o achieves comparatively better performance than Qwen-2.5 and DeepSeek-R1. Interestingly, we found that the ensemble technique cannot further improve the performance of LLM-based classifiers.
@@
In the future, we plan to conduct a more in-depth investigation of how practitioners deal with architecture violations in code review, what kind of practices and tools they use in handling architecture violations, and evaluate whether our approach can support the identification of architecture violations in their daily job (e.g., through controlled experiments to compare the effectiveness of our approach with human-centric practices). Moreover, we want to evolve the classifiers for identifying fine-grained types of architecture violations based on embeddings for code review comments and a larger dataset of more projects from both open-source and industry projects. In particular, a bigger picture is to delve into a more in-depth investigation of the performance of trained classifiers as plug-ins or tools using more real-world scenarios. The goal of such plug-ins or tools for code review platforms is to provide warnings of potential architecture violations to practitioners (e.g., developers and reviewers) for them to re-check the related code changes during code review.
Data Availability
The replication package of this work has been made available at (onlinepackage, ).